# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v2

@@ 5.4 Calibration
The forward and call prices are computed using (4.14) and (4.15), respectively, where the true price uses the true parameters $\bf b$}, $\boldsymbol{\vartheta}$ and the predicted price uses the estimated parameters $\widehat{\bf b}$, $\widehat{\boldsymbol{\vartheta}}$. To evaluate (4.15), we use numerical integration rather than FFT as done by Carr and Madan. A discussion on the improved accuracy and speed of this is given in (pasc, , Section 15.2.3).
@@
Our approach estimates $\boldsymbol{\vartheta}$ under $\mathbb{P}$ and calibrates $\bf h$} under $\mathbb{Q}_{{\bf h}}$, in contrast to the approach using the mean-correcting martingale measure, such as in Gardini and Sabino gs22 . The latter requires the model be specified under a risk-neutral measure $\mathbb{Q}$ rather than $\mathbb{P}$, and then calibrated to forward and call prices. An issue then arises with the estimation of joint parameters, such as those affecting the correlation, which should be calibrated to the prices of other bivariate energy derivatives, but these may not be readily traded in the market, while estimating them using historical correlations under $\mathbb{P}$ is inconsistent with the model. In contrast, our method combines both the observed energy prices $\bf S$} under $\mathbb{P}$ and univariate energy derivative prices to fit the model in a theoretically sound way. In fact, it is unclear whether our model can be estimated only by calibrating to derivative prices, as it may be the case that $\boldsymbol{\vartheta}$ and $\bf h$} would become unidentifiable. This issue arises if energy prices follow the simpler Black-Scholes model.
@@ 5.5 Pricing
    Adjust the $\bf x$}-grid so that $\log({\bf S}(t)/K)$ is on it. Do this by setting $\bf j$$^{*}$ as the index of the grid point nearest $\log({\bf S}(t)/K)$, and then adjust the values of $\delta_{\theta_{1}},\delta_{\theta_{2}}$ such that $\log({\bf S}(t)/K)=[{\bf x}]_{{\bf j}^{*}}$ holds.
@@
Note that the FFT method for spread option pricing is different from the Fourier inversion method for estimation. The former uses the discrete inverse Fourier transform, while the latter uses the discrete Fourier transform.
$\Box$
@@
6 Numerical results
We now present numerical results on option pricing in our model based on the implementation details in Section 5. Results in Sections 6.16.2 are under the true model. Results in Sections 6.36.4 are from a simulation study using fitted models.
@@ 6.1 Spread option pricing under true parameters
Using the true parameters, Table 1 gives the spread option price using the FFT method with $N=2^{11}$ grid points along each axis and $\overline{\theta}=80$, and the Monte Carlo method with $N=10^{6}$ simulations and $\widetilde{n}=10^{5}$ subintervals.
@@
$\text{Average error}=\frac{1}{q}\sum_{i=1}^{q}|O_{i}-E_{i}|,$
where $O_{i}$ is the true price of the spread option with strike price $K_{i}$, and $E_{i}$ is the price of the spread option computed in another way. We take the true price to be from the FFT method with $N=2^{11}$, $\bar{u}=80$, and it is compared with the prices using the discretization parameters $N=2^{k}$, $k=6,7,\dots 10$, $\overline{\theta}=40,80$. Figure 2 shows how this affects the average error of the FFT method.
@@ 6.2 Spread option parameter sensitivities
Figure 3 shows the effect of the LDOUP parameters on the spread option price with strike price $K=3.6$ and all others held constant. The parameters $a,\alpha_{1},\alpha_{2},\Sigma_{12}$ are plotted over their full range subject to the parameter constraints, the same is also true for the lower limit of $\Sigma_{11},\Sigma_{22}$.
The spread option price decreases as the rate of mean reversion $\lambda$ increases, as this reduces variability. The plots for the kurtosis parameters $\alpha_{1}$ and $\alpha_{2}$ have a point of nondifferentiability at $\alpha_{2}=\alpha_{1}$ due to
$\operatorname{Cov}_{\mathbb{Q}_{{\bf h}}}(X_{1}(T_{o}),X_{2}(T_{o}){\,|\,}{\cal F}_{T_{e}})=\frac{1}{2}{(1-e^{-2\lambda\tau})}a((\alpha_{1}\wedge\alpha_{2})\Sigma_{12}+\alpha_{1}\alpha_{2}\mu_{1}\mu_{2}).$
As $\alpha_{2}<\alpha_{1}$ decreases, the covariance decreases causing the price to increase, while as $\alpha_{2}>\alpha_{1}$ increases, the covariance increases but the slope is near zero because $\mu_{1}\mu_{2}$ is as well. Some parameters may have surprising effects given that they affect moments other than what is considered their primary one. For example, it may be expected that increasing the skewness parameters $\mu_{1}$ and $\mu_{2}$ have opposite effects, but actually both tend to increase the variance of the spread under $\mathbb{Q}_{{\bf h}}$ in asymmetric ways, which causes the price to increase. As $\Sigma_{22}$ increases, this increases the variance of the spread because $\Sigma_{11},\Sigma_{12}$ are held constant, but $\rho$ is not, and hence the price increases.
@@ 6.3 Fitting the model
$\widetilde{\mu}_{k}:={}$ $\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[R_{k}(T_{e},T_{o}){\,|\,}{\cal F}_{T_{e}}]$
@@
Table 2 gives the true value, mean estimate, estimated standard error, and estimated relative root mean square error (RRMSE) of the seasonal parameters $\bf b$}, LDOUP parameters $\boldsymbol{\vartheta}$, MPRs ${\bf h}^{f}=(h_{1}^{f},h_{2}^{f})$ and ${\bf h}^{c}=(h_{1}^{c},h_{2}^{c})$ from calibrating to forward and call prices, respectively, and their corresponding risk-neutral drifts $\widetilde{\boldsymbol{\mu}}^{f}=(\widetilde{\mu}_{1}^{f},\widetilde{\mu}_{2}^{f})$, $\widetilde{\boldsymbol{\mu}}^{c}=(\widetilde{\mu}_{1}^{c},\widetilde{\mu}_{2}^{c})$
@@
From these results, most model parameters are estimated with moderately low RRMSE of approximately 10–30%, other than those with true values near 0. Related to this, note that errors in estimating the model parameters are compensated for by the calibration. For instance, $h_{1}^{c}$ and $h_{2}^{c}$ have RRMSEs of 589% and 2612%, which can take them far from their true value, so that $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ can have small RRMSEs of 49% and 12%, close to their true value. In fact, on an absolute scale as seen by the vertical line in Figure 5, the differences between the true and estimated values of $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ are negligible. Ultimately, it is $\widetilde{\boldsymbol{\mu}}$ that has a first-order effect on the energy derivative prices, and the fitted models should primarily be judged on how accurately they prices spread options.
Figure 6: The true call prices on $S_{1}$ (black) over various strike prices $K_{i}$, and the predicted call price using the true MPR $\bf h$} (green) and calibrated MPR $\widehat{\bf h}^{c}$ (red).
@@ 6.4 Spread option pricing under the fitted model
Finally, over the 1000 simulations, the results from using the fitted model to price spread options are presented in Table 3, with the calibration to forward and call prices shown separately. The spread option prices are computed by the FFT method with $N=2^{8}$, $\bar{u}=40$, which gives sufficient accuracy as noted in Section 6.1. The mean prices and their RRMSEs are given.
Calibrated to forwards Calibrated to calls $K_{i}$ True price Mean fitted price RRMSE (%) Mean fitted price RRMSE (%) 0.4 9.31079 8.34185 28.87673 8.14641 16.05140 1.2 9.03506 8.07630 29.19487 7.88303 16.37409 2.0 8.76650 7.81817 29.51487 7.62714 16.69879 2.8 8.50502 7.56764 29.83218 7.37888 17.02162 3.6 8.25053 7.32462 30.14680 7.13819 17.34162 4.4 8.00290 7.08870 30.46282 6.90463 17.66241 5.2 7.76200 6.85969 30.77809 6.67802 17.98405 6.0 7.52774 6.63787 31.08973 6.45865 18.30082 6.8 7.29998 6.42256 31.40161 6.24581 18.61972 7.6 7.07862 6.21396 31.71286 6.03969 18.93635 8.4 6.86350 6.01140 32.02554 5.83965 19.25737 9.2 6.65454 5.81541 32.33447 5.64619 19.57346 10.0 6.45158 5.62567 32.64130 5.45901 19.88679
@@
The results show that the spread options are priced less accurately when calibrated to forwards, with an average RRMSE of 30.8%, than to calls, where it is 18.0%. This shows how errors in fitting the model, from 2001 observations and 40 calibration instruments, propagate to pricing errors. As the sample size increases, so that the model parameters are estimated more accurately, the pricing error should approach 0.

### v3

@@ 5.1.4 Calibration
The forward and call prices are computed using (4.13) and (4.16), respectively, where the true price uses the true parameters $\bf b$}, $\boldsymbol{\vartheta}$ and the predicted price uses the estimated parameters $\widehat{\bf b}$, $\widehat{\boldsymbol{\vartheta}}$. To evaluate (4.16), we use numerical integration rather than FFT as done by CM99 . A discussion on the improved accuracy and speed of this is given in (pasc, , Section 15.2.3).
@@
Our approach estimates $\boldsymbol{\vartheta}$ under $\mathbb{P}$ and calibrates $\bf h$} under $\mathbb{Q}_{{\bf h}}$, in contrast to the approach using the mean-correcting martingale measure, such as in Gardini and Sabino gs22 . The latter requires the model be specified under a risk-neutral measure $\mathbb{Q}$ rather than $\mathbb{P}$, and then calibrated to forward or call prices. An issue then arises with the estimation of joint parameters, such as those affecting the correlation, which should be calibrated to the prices of other multi-asset energy derivatives, but these may not be liquidly traded in the market. To address this, one way is to calibrate the joint parameters such that the historical correlation of the log returns of $\bf S$} under $\mathbb{P}$ matches that of the model correlation under $\mathbb{Q}$ as a proxy. This effectively assumes that the $\mathbb{P}$ and $\mathbb{Q}$ correlations of the log return are equal, which is inconsistent with our model. More generally, one should expect that any change of measure can potentially change the correlation. An additional issue of this being an underdetermined problem, as there are 2 joint parameters but 1 correlation constraint, leaving 1 degree of freedom, is raised in (gs22, , Footnote 2). We avoid this method.
In contrast, our method combines both the observed energy prices $\bf S$} under $\mathbb{P}$ and univariate energy derivative prices to fit the model in a theoretically sound way. In fact, it is unclear whether our model can be estimated only by calibrating to derivative prices, as it may be the case that $\boldsymbol{\vartheta}$ and $\bf h$} would become unidentifiable. This issue arises if energy prices follow the simpler Black-Scholes model.
@@ 5.1.5 Pricing
    Adjust the $\bf x$}-grid so that $\log({\bf S}(t)/K)$ is on it. Do this by setting $\bf j$$^{*}$ as the index of the grid point nearest $\log({\bf S}(t)/K)$, and then adjust the values of $\delta_{\theta_{1}},\delta_{\theta_{2}}$ such that $\log({\bf S}(t)/K)=[{\bf x}]_{{\bf j}^{*}}$ holds, provided all elements of ${\bf x}]_{{\bf j}^{*}}$ are nonzero.
@@
Note that the FFT used for spread option pricing is different from the one used in the Fourier inversion method for estimation. The former uses the discrete inverse Fourier transform, while the latter uses the discrete Fourier transform.   $\Box$
@@
5.2 Real data example
In the real data example, we fit a model to Australian electricity and base load futures prices and then price spread options under the model. Our implementation mostly follows Section 5.1, except for the differences outlined here. All unmentioned details remain the same as before.
The data consists of 8 years of daily average electricity prices for the Australian states NSW and Victoria in the NEM sourced from AEMO (Australian Energy Market Operator). These prices were obtained from 2018-01-05 to 2026-01-05. This is used to estimate the model parameters. The electricity price data consists of 2923 daily observations, so that $m=2922$ and $\Delta=\frac{1}{365.25}$ accounting for leap years.
Let ${\bf S}^{\prime}=(S^{\prime}_{1},S^{\prime}_{2})$, where $S^{\prime}_{1}(t)$ and $S^{\prime}_{2}(t)$ are the NSW and Victoria daily average electricity at time $t$, respectively. The observations of $S_{1}^{\prime}$ are always positive, but 5.9% of the observations of $S_{2}^{\prime}$ are negative with a minimum price $\mathdollar 73.02$, so a geometric model cannot be used to model $S_{2}^{\prime}$ directly. Using Remark 4.12 to deal with this, we let $S_{1}(t)=S^{\prime}_{2}(t)+c$, $S_{2}(t)=S_{1}^{\prime}(t)$, $t\geq 0$, and fit the geometric model to ${\bf S}=(S_{1},S_{2})$, where $c$ is a constant that makes the observations of $S_{1}$ positive. The NEM has a negative price floor, so such a $c$ exists. We take $c=100$, however, different choices lead to different models.
We assume the seasonality function is
$\Lambda_{k}(t)={}$ $b_{0k}+b_{1k}t+b_{2k}\cos\mathopen{}\left(\frac{2\pi}{p_{1}}t\right)\mathclose{}+b_{3k}\sin\mathopen{}\left(\frac{2\pi}{p_{1}}t\right)\mathclose{}$
$+b_{4k}\cos\mathopen{}\left(\frac{2\pi}{p_{2}}t\right)\mathclose{}+b_{5k}\sin\mathopen{}\left(\frac{2\pi}{p_{2}}t\right)\mathclose{},\quad t\in[0,T],$
with periods $p_{1}=1$ and $p_{2}=7\Delta$, which correspond to yearly and weekly cyclical components, respectively. The inclusion of weekly seasonality in electricity prices makes the model more realistic.
The risk-free rate is assumed to be $r=0.03715$. This only affects the discount term for the pricing of spread options and plays no role in fitting the model.
At the current time $T_{e}=8$, we price spread options on the original prices $(S_{1}^{\prime},S_{2}^{\prime})$ whose payoff is given by (4.20), where the strike prices are $K^{\prime}_{i}$, $i=1,\dots,13$, which are equally spaced values from 20 to 80, and the time to maturity is $\tau=0.5$. This is done by computing the price of spread options on $(S_{1},S_{2})$ using (4.21).
The data also consists of base load calendar quarter electricity futures prices for NSW and Victoria with 8 contract periods Q1 2026 to Q4 2027 on the day 2026-01-05 sourced from ASX Energy (Australian Stock Exchange). This is used to calibrate the model, so there are $q=16$ calibration instruments. The base load futures prices on $S_{k}$, which is $\overline{F}_{k}(t,T_{1},T_{2})$, is computed using (4.22) and (4.14) for $k=1,2$. For example, since the current time is 2026-01-05, the Q1 2026 base load futures price has $t=T_{e}+5\Delta$, $T_{1}=T_{e}$, $T_{2}=T_{e}+90\Delta$, and is computed using 5 observed electricity prices from 2026-01-01 to 2026-01-05 by (4.15).
In the FFT method, we set $N=2^{8}$, $\bar{\theta}=40$, and the grid is always adjusted so that $K=50$ is on it.
6 Simulated data results
We now present results from simulated data examples based on the implementation details in Section 5.1. Results in Sections 6.16.2 are under the true model. Results in Sections 6.36.4 are from a simulation study using fitted models.
@@ 6.1 Spread option pricing under true parameters
Using the true parameters, Table 1 gives the spread option price using the FFT method with $N=2^{11}$ grid points along each axis and $\overline{\theta}=80$, and the Monte Carlo method with $N=10^{6}$ simulations and $\widetilde{n}=10^{5}$ subintervals by discretizing the SDE.
@@
$\text{Average error}=\frac{1}{q}\sum_{i=1}^{q}|O_{i}-E_{i}|,$ (6.1)
where $O_{i}$ is the true price of the spread option with strike price $K_{i}$, and $E_{i}$ is the price of the spread option computed in another way. We take the true price to be from the FFT method with $N=2^{11}$, $\bar{\theta}=80$, and it is compared with the prices using the discretization parameters $N=2^{k}$, $k=6,7,\dots 10$, $\overline{\theta}=40,80$. Figure 2 shows how this affects the average error of the FFT method.
@@ 6.2 Spread option parameter sensitivities
Figure 3 shows the effect of the LDOUP parameters on the spread option price with strike price $K=3.6$ and all other parameters held constant. We refer to (MiSz17, , Remark 4) for a discussion about the interpretation of the parameters of the WVAG process. The parameters $a,\alpha_{1},\alpha_{2},\Sigma_{12}$ are plotted over their full range subject to the parameter constraints, the same is also true for the lower limit of $\Sigma_{11},\Sigma_{22}$.
The spread option price decreases as the rate of mean reversion $\lambda$ increases, as this reduces the variance of the spread. The plots for the kurtosis parameters $\alpha_{1}$ and $\alpha_{2}$ have a point of nondifferentiability at $\alpha_{2}=\alpha_{1}$ due to the minimum function occurring in various terms of
$\operatorname{Cov}_{\mathbb{Q}_{{\bf h}}}(X_{1}(T_{o}),X_{2}(T_{o}){\,|\,}{\cal F}_{T_{e}})=\frac{1}{2}{(1-e^{-2\lambda\tau})}\mathopen{}\left({[{\boldsymbol{\Sigma}}_{{\bf h}}]}_{12}+\frac{{[\boldsymbol{\mu}_{{\bf h}}]}_{1}{[\boldsymbol{\mu}_{{\bf h}}]}_{2}}{a}\right)\mathclose{},$
which follows from Corollary 4.3, and where for example, ${[{\boldsymbol{\Sigma}}_{{\bf h}}]}_{12}=a\Sigma_{12}(\alpha_{1}\wedge\alpha_{2})/K_{{\bf h}}$. This covariance in turn affects the variance of the spread under $\mathbb{Q}_{{\bf h}}$. Some parameters may have surprising effects given that they affect moments other than what is considered their primary one. For example, it may be expected that increasing the skewness parameters $\mu_{1}$ and $\mu_{2}$ have opposite effects, but actually both tend to increase the variance of the spread under $\mathbb{Q}_{{\bf h}}$ in asymmetric ways, which causes the price to increase. As variance parameter $\Sigma_{22}$ increases, this increases the variance of the spread because $\Sigma_{11},\Sigma_{12}$ are held constant, but the correlation ${\Sigma_{12}}/\sqrt{\Sigma_{11}\Sigma_{22}}$ is not, and hence the price increases.
@@ 6.3 Fitting the model
$\widetilde{\mu}_{k}:={}$ $\mathbb{E}_{\mathbb{Q}_{{\bf h}}}[R_{k}(T_{e},T_{o}){\,|\,}{\cal F}_{T_{e}}]$ (6.2)
@@
Table 2 gives the true value, mean estimate, estimated standard error, and estimated relative root mean square error (RRMSE) of the seasonal parameters $\bf b$}, LDOUP parameters $\boldsymbol{\vartheta}$, MPRs ${\bf h}^{f}=(h_{1}^{f},h_{2}^{f})$ and ${\bf h}^{c}=(h_{1}^{c},h_{2}^{c})$ from calibrating to forward and call prices, respectively, and their corresponding risk-neutral drifts $\widetilde{\boldsymbol{\mu}}^{f}=(\widetilde{\mu}_{1}^{f},\widetilde{\mu}_{2}^{f})$, $\widetilde{\boldsymbol{\mu}}^{c}=(\widetilde{\mu}_{1}^{c},\widetilde{\mu}_{2}^{c})$.
@@
From these results, most model parameters are estimated with moderately low RRMSE of approximately 10–30%, other than those with true values near 0. Related to this, note that errors in estimating the model parameters are compensated for by the calibration. For instance, $h_{1}^{c}$ and $h_{2}^{c}$ have RRMSEs of 589% and 2612%, which can take them far from their true value, so that $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ can have small RRMSEs of 49% and 12%, close to their true value. In fact, on an absolute scale as seen by the vertical line in Figure 5, the differences between the true and estimated values of $\widetilde{\mu}^{c}_{1}$ and $\widetilde{\mu}^{c}_{2}$ are negligible. Ultimately, it is $\widetilde{\boldsymbol{\mu}}$ that has a first-order effect on the energy derivative prices, and the fitted models should primarily be judged on how accurately they price spread options.
Figure 6: The true call prices on $S_{1}$ (black) over various strike prices $K_{i}$, and the predicted call prices using the true MPR $\bf h$} (green) and calibrated MPR $\widehat{\bf h}^{c}$ (red).
@@ 6.4 Spread option pricing under the fitted model
Finally, over the 1000 simulations, the results from using the fitted model to price spread options are presented in Table 3, with the calibration to forward and call prices shown separately. The spread option prices are computed by the FFT method with $N=2^{8}$, $\bar{u}=40$, which gives sufficient accuracy as noted in Section 6.1. The true prices are from Table 1, computed with $N=2^{11}$, $\bar{u}=80$. The mean prices and their RRMSEs are given.
Calibrated to forwards Calibrated to calls $K_{i}$ True price Mean fitted price RRMSE (%) Mean fitted price RRMSE (%) 0.4 9.31079 8.28003 29.28306 8.14641 16.05140 1.2 9.03505 8.01567 29.60813 7.88303 16.37409 2.0 8.76650 7.75871 29.93505 7.62714 16.69879 2.8 8.50503 7.50936 30.25915 7.37888 17.02162 3.6 8.25053 7.26751 30.58043 7.13819 17.34162 4.4 8.00289 7.03275 30.90307 6.90463 17.66241 5.2 7.76200 6.80489 31.22491 6.67802 17.98405 6.0 7.52773 6.58422 31.54294 6.45865 18.30082 6.8 7.29998 6.37004 31.86120 6.24581 18.61972 7.6 7.07861 6.16257 32.17869 6.03969 18.93635 8.4 6.86350 5.96113 32.49764 5.83965 19.25737 9.2 6.65454 5.76624 32.81268 5.64619 19.57346 10.0 6.45158 5.57759 33.12549 5.45901 19.88679
@@
The results show that the spread options are priced less accurately when calibrated to forwards, with an average RRMSE of 31.2%, than to calls, where it is 18.0%. This shows how errors in fitting the model, from 2001 observations and 40 calibration instruments, propagate to pricing errors. As the sample size increases, so that the model parameters are estimated more accurately, the pricing error should approach 0.
@@
7 Real data results
We present results from a real data example based on the implementation details in Section 5.2, where we fit the model to Australian electricity and base load futures prices, and price spread options.
Figure 7 plots the sample path of the log price $Y_{k}(t)$ and the estimated seasonality function $\widehat{\Lambda}_{k}(t)$, in which a small weekly cyclical component is visible. Table 4 gives the estimated parameters of the fitted model, where $(\widetilde{\mu}_{1},\widetilde{\mu}_{2})$ are the risk-neutral drifts defined in (6.2).
Figure 7: Sample path of the log price $Y_{k}(t)$ (black) and the estimated seasonality function $\widehat{\Lambda}_{k}(t)$ (red) for $k=1,2$.
Parameter Estimate
$b_{01}$ $5.19036$
$b_{11}$ 0.01937-0.01937
$b_{21}$ 0.14974-0.14974
$b_{31}$ $0.07126$
$b_{41}$ 0.01048-0.01048
$b_{51}$ 0.06942-0.06942
$b_{02}$ $4.22798$
$b_{12}$ $0.04056$
$b_{22}$ 0.19417-0.19417
$b_{32}$ $0.08653$
$b_{42}$ 0.01921-0.01921
$b_{52}$ 0.13987-0.13987
Parameter Estimate
$\lambda$ $123.14798$
$a$ $2.15520$
$\alpha_{1}$ $0.34922$
$\alpha_{2}$ $0.31290$
$\mu_{1}$ 0.01398-0.01398
$\mu_{2}$ $0.13247$
$\Sigma_{11}$ $0.20842$
$\Sigma_{22}$ $0.49101$
$\Sigma_{12}$ $0.26145$
$h_{1}$ $0.80684$
$h_{2}$ 0.48216-0.48216
$\widetilde{\mu}_{1}$ $0.44854$
$\widetilde{\mu}_{2}$ $0.78568$
Table 4: The estimate of the seasonal parameters, LDOUP parameters, MPRs, risk-neutral drifts.
Figure 8: Kernel density estimate (black) and fitted (red) pdf of $Z_{1}^{*}(\Delta)$ (left) and $Z_{2}^{*}(\Delta)$ (right) under $\mathbb{P}$ on the original scale (top) and log scale (bottom).
To assess the goodness-of-fit, we plot in Figure 8 the pdf of the marginal components of the innovation term ${\bf Z}^{*}(\Delta)$ under $\mathbb{P}$ obtained by the kernel density estimate and the fitted distribution. The observations of ${\bf Z}^{*}(\Delta)$ were obtained by (5.6) and (2.10) with the estimate of $\lambda$. The plot effectively compares the empirical and fitted distribution of the residual for each marginal component on the original scale and log scale. The fit appears to be reasonably good. As the MLE is not a robust estimation method, it has fit the tails well on the log scale, but at the expense of the fit in the center as shown on the original scale. In fact, if not using the MLE, it is possible to obtain a visually near-perfect fit on the original scale but a worse fit to the tails on the log scale. To assess the joint distribution, the theoretical and sample value of the correlation $\operatorname{Corr}({\bf Z}_{1}^{*}(\Delta),{\bf Z}_{2}^{*}(\Delta))$ are 0.55 and 0.47, respectively, which is reasonably close.
Another way to assess goodness-of-fit is to compare the predicted and observed prices of the base load futures prices for each of the 8 contract periods on the day 2026-01-05 as shown in Figure 9. Note that from the base load futures prices on $S_{1}$, subtracting $c$ gives the base load futures prices for Victoria. The average percentage error, which is defined by (6.1) with summand $|O_{i}-E_{i}|/E_{i}$ instead, is 4.62%. In addition, the predicted and observed futures curves have a qualitatively similar shape. Overall, the results suggest the model is a reasonable fit to the data despite some imperfections.
Figure 9: The observed base load futures price (black) on $S_{1}$ and $S_{2}$ for the contract periods Q1 2026 to Q4 2027, and the predicted futures prices (red) from the fitted model.
The results here can also be compared to those in the literature. In bsb04 , a univariate LDOUP with a NIG BLDP was fit separately to electricity and gas price data, and their Figure 5 and 6 is similar to Figure 8 here. They make the same finding regarding the fit to the center and tail. In cf05 , a jump-diffusion BDLP is used instead, and their Figure 9 has a slightly worse fit compared to Figure 9 here. In contrast to both these univariate models, we have fit a multivariate model.
Finally, using the fitted model on the day 2026-01-05, we give the price of the spread option on $(S_{1},S_{2})$ with strike prices $K_{i}$, and the corresponding price of the spread option on the original prices $(S_{1}^{\prime},S_{2}^{\prime})$ with strike prices $K^{\prime}_{i}$ in Table 5.
$K_{i}$ Price on $(S_{1},S_{2})$ $K^{\prime}_{i}$ Price on $(S^{\prime}_{1},S^{\prime}_{2})$
20 66.32141 80 6.27083
25 62.08124 75 6.93864
30 57.93256 70 7.69793
35 53.88806 65 8.56142
40 49.96123 60 9.54256
45 46.16612 55 10.65543
50 42.51608 50 11.91336
55 39.02593 45 13.33119
60 35.70724 40 14.92047
65 32.57084 35 16.69204
70 29.62514 30 18.65432
75 26.87572 25 20.81287
80 24.32528 20 23.17042
Table 5: Spread option prices on $(S_{1},S_{2})$ for the fitted models over various strike prices $K_{i}$, and the corresponding spread option prices on $(S^{\prime}_{1},S^{\prime}_{2})$ over the strike prices $K_{i}^{\prime}$.